OpenStudy (liizzyliizz):

im graphing functions that i need to find the domain and range of, ( I know what domain and range is) but is there a quick way to figure it out and writing it correctly in interval notation?) I think im over thinking it, but i want to make sure im doing my work correctly...

OpenStudy (anonymous):

where's your work

OpenStudy (liizzyliizz):

ill upload the file hold on.

OpenStudy (liizzyliizz):

its basically like graphs like \[y=\sqrt{x} , y=\sqrt[3]{x?} , y= e ^{x} , y=lnx , y=2^5 , y= 1/x , y= x^2-4 , y= x^2+4x+3, y=sinx , y= \sqrt{x-2}, y=\sqrt{4-x^2 , and y=|x+3|-2 }\]

OpenStudy (anonymous):

? are you trying to say that x = your domain and y = your range because you don't write it that way

OpenStudy (anonymous):

2aroub, u been drinking...?

OpenStudy (liizzyliizz):

its basically like graphs like \[y=\sqrt{x} , y=\sqrt[3]{x?} , y= e ^{x} , y=lnx , y=2^5 , y= 1/x , y= x^2-4 , y= x^2+4x+3, y=sinx , y= \sqrt{x-2}, y=\sqrt{4-x^2 , and y=|x+3|-2 }\]

OpenStudy (liizzyliizz):

ahh i dont think you can see them all, but basically its like im not sure if im calculating them right, and if i can do one correctly i can probbaly figure out the rest on my own.

OpenStudy (anonymous):

well what did youdo for y = sqroot(x)

OpenStudy (liizzyliizz):

i wrote that x is greater than or equal to 0 and wrote [0, the infinity sign)

OpenStudy (anonymous):

The short answer is u have to figure it out, there is no quick way...

OpenStudy (liizzyliizz):

i know the rules for some of them but its like im kind of stuck on how to figure out the others..

OpenStudy (liizzyliizz):

especially the range, i know its pretty simple but its been a while since ive done these problems and im probably just over thinking but i dont want to get them all wrong because of a simple mistake...

OpenStudy (anonymous):

graphing wold be the easiest

OpenStudy (liizzyliizz):

i graphed them, and im looking at them, but i feel stuck, and its frustrating because i dont know why i cant figure it out, (i never liked graphs)

OpenStudy (anonymous):

your domain is going to be what values can x be and range will be all y values

OpenStudy (anonymous):

See, after a while u just get to know most of them.... Usually its just R in and out but of course they are putting all these ones to get u to learn the odd ones...

OpenStudy (anonymous):

so your domain for like sq root is usually going to be 0,infinity

OpenStudy (anonymous):

U need to know what the graphs of things line ln x and e^x look like without having to graph them...

OpenStudy (liizzyliizz):

I understand that, but for example \[y=\sqrt[3]{x?} \] would the domain be \[\left( -\infty, \infty \right)\]? but what would the range be , the same?

OpenStudy (anonymous):

cube root -8?

OpenStudy (liizzyliizz):

lol ok i just sounded stupid, how would you right the range then.

OpenStudy (anonymous):

R

OpenStudy (liizzyliizz):

write* eep

OpenStudy (anonymous):

Estudier might know but i'm not 100 % sure i know you could find it by calculus but as for now i guess all you can assume is that since domain can only be positive that the numbers generate will be positive to infinity

OpenStudy (anonymous):

Domain for cube root is R as well.

OpenStudy (anonymous):

Positive is just for square root under radical

OpenStudy (anonymous):

so you could like evaluate to show how you found the rane by like saying f(0) = ... and then f(infinity)=

OpenStudy (liizzyliizz):

so \[\left[ 0,\infty \right]\] for both?

OpenStudy (anonymous):

No need to evaluate, just state that domain and range is R.

OpenStudy (liizzyliizz):

i feel like i understand what you are trying to say but its frustrating lol

OpenStudy (anonymous):

What's this obsession with infinity?

OpenStudy (anonymous):

yeah you could just say but my teacher would be like "HOW DID YOU COME TO THE ASSUMPTION, YOU DIDNT SHOW,WRONG"

OpenStudy (anonymous):

because your values can go to infinity with those functions

OpenStudy (anonymous):

idk why the hes says the domain can be all real numbers for square roots

OpenStudy (liizzyliizz):

i feel like in pre cal i used infinity a lot, and my teacher would include it alot in interval notation, now that im doing work for Ap calc, i feel nervous im trying to refreshen my memory, and like outkast said alot of these functions can go to infinity so i just assumed i would need it in these cases.. but im not sure which is why im asking

OpenStudy (anonymous):

idk why the hes says the domain can be all real numbers for square roots We have been talking about cube roots...

OpenStudy (anonymous):

U know the domain and the range are sets, right?

OpenStudy (liizzyliizz):

that is just one example... i have 12 to do in total lol. im trying to figure out how to do a feel to continue with the rest.

OpenStudy (liizzyliizz):

and yes.

OpenStudy (anonymous):

The set R is an infinite set.

OpenStudy (anonymous):

ie all real numbers....

OpenStudy (anonymous):

And u know infinity is not a number, right?

OpenStudy (liizzyliizz):

yess.... im not that slow -_- im just trying to refreshen my memory on this im not an idiot...

OpenStudy (anonymous):

I'm not entirely talking to u (and I am not saying u r slow). There seems to be some misunderstandings here, it would be good to clear them up...

OpenStudy (liizzyliizz):

yeah it would be good lol. ugh im just frustrated with these graphs

OpenStudy (anonymous):

Well, let's go 1 by 1... y = sqrt x was first, was it?

OpenStudy (liizzyliizz):

yes

OpenStudy (anonymous):

U got that one or not?

OpenStudy (liizzyliizz):

i know it cant be negative numbers.. i wrote 0,positive infinity for both...

OpenStudy (anonymous):

Personally I would say x>=0 but if that is what u have been taught to put....

OpenStudy (anonymous):

And because of that y is also >=0

OpenStudy (anonymous):

Because I don't understand what sqrt(infinity) means...

OpenStudy (liizzyliizz):

thats what i wrote i didnt know how to put it in interval

OpenStudy (liizzyliizz):

not sqrt infinity, but in these cases the numbers can be endless the line goes on thats what they mean

OpenStudy (anonymous):

U can write [0, infinity) as an "interval" if u want to. To me it makes no sense as a domain because u cannot take the sqrt of infinity.

OpenStudy (liizzyliizz):

ur not taking the sqrt of infinity it just means the lines goes on it could be any number going on

OpenStudy (anonymous):

I understand what u r saying but the domain is a set, the set is called R. U are writing an interval, which is fine, but then saying that is the domain which is where u lose me. Look here http://en.wikipedia.org/wiki/Domain_of_a_function

OpenStudy (anonymous):

Not R in this case, 0 or positive integers, sorry...t

OpenStudy (liizzyliizz):

i see what u are saying.

OpenStudy (anonymous):

The set of positive integers plus zero I would usually call N

OpenStudy (anonymous):

The problem here maybe that I don't really know what it is u have been taught up to now......

OpenStudy (anonymous):

Have u encountered these terms like R, Z, N before..?

OpenStudy (liizzyliizz):

I remember being taught this, but that was a while back, now im preparing for AP calc and im looking at this in terror lol. i did, but briefly

OpenStudy (anonymous):

U said u knew that domain and range is a set, so I just assumed..

OpenStudy (anonymous):

What is AP Calc?

OpenStudy (liizzyliizz):

Advanced placement calculus AB?

OpenStudy (anonymous):

placement to uni?

OpenStudy (liizzyliizz):

yeah,

OpenStudy (anonymous):

Then I think u should go with the correct way and describe the domain and range as sets....

OpenStudy (liizzyliizz):

hmm, i shall keep looking the domain and range is the only thing i left blank :/

OpenStudy (anonymous):

Do u know how to use Wolfram?

OpenStudy (liizzyliizz):

no

OpenStudy (anonymous):

It will make your life easier for the graphs...

OpenStudy (anonymous):

Or simpler http://www.wolframalpha.com/input/?i=y%3De^x

OpenStudy (anonymous):

U just type in the function and it will give u the graph and a bunch other stuff...

OpenStudy (liizzyliizz):

thank you !

OpenStudy (anonymous):

Last thing, u can specify the domain and range of plot like this (try with some numbers of your own).

OpenStudy (anonymous):

Careful with the cube root, doesn't quite work in the way u would normally expect

OpenStudy (anonymous):

U have to do this way (Re means reals) http://www.wolframalpha.com/input/?i=Re+y%3D+x^3+

OpenStudy (anonymous):

Anyway, hope all goes well for u...

OpenStudy (anonymous):

It seems that I have misled u somewhat, u are apparently permitted to describe a function range and domain simply using interval notation with (-inf, inf) being R. Seems to me u should not get into trouble the other way, however, since that is also correct.